OpenStudy (anonymous):

10. n and p are integers greater than 1 5n is the square of a number 75np is the cube of a number. The smallest value for n + p is...

OpenStudy (anonymous):

don't you have a textbook from class, and didin't your teacher explain this?

OpenStudy (anonymous):

I had a textbook hmm about 15 years ago when I was in a class.. and Im sure my teacher did

OpenStudy (anonymous):

i think n = 5 and p = 9 n+ p = 14

OpenStudy (anonymous):

you're correct.. in fact I have the answer, but Im not sure how it was obtained.. I'd like to see how.

OpenStudy (anonymous):

well n= 5 gives 5n =25 then 75n = 75 * 5 = 375 and i guessed 9 375 * 9 = 3375 - taking the cubic root of this gives 15.

OpenStudy (anonymous):

so basically its taking educated guesses to arrive at the answer?

OpenStudy (anonymous):

the second part was a guess - there's probably a more logical way to do this

OpenStudy (anonymous):

the first part had to be n = 5 if you want the lowest possible value

OpenStudy (anonymous):

This is actually a question on a practice GRE exam.. I can't imagine solving that logically in under 1 minute (about the time you have per question on the quantitative section) I'll post the exam guide's explanation.. I dont really get it.. But you probably will

OpenStudy (anonymous):

Explanation: The smallest value for n such that 5n is a square is 5. 75np can now be written as 75 x 5 x p. This gives prime factors.... 3 x 5 x 5 x 5 x p To make the expression a perfect cube, p will have to have factors 3 x 3 , and hence p =9 n + p = 5 + 9 = 14

OpenStudy (anonymous):

yes! - i should have thought of that way- if a number two duplicate primes as factors and three triplicates it is a perfect cube

OpenStudy (anonymous):

I'll have to look up perfect cubes and prime factors I guess.. I'm just not getting this one.. Really rusty with my math..

OpenStudy (anonymous):

Thanks though

OpenStudy (anonymous):

if a number has even quantities of prime factors it is a perfect square so for example to make 2 x 2 x 3 (12) a perfect square you multiply by 3 = 2 x 2 x 3 x 3 = 36

OpenStudy (anonymous):

ah... ok.. I get that..